OpenStudy (anonymous):

Solve the following equation on the interval [0,2pi). sin x-2 sin x cos x = 0

OpenStudy (xapproachesinfinity):

how about factoring that sinx what do get if you did

OpenStudy (anonymous):

1-cos x?

OpenStudy (xapproachesinfinity):

where did sinx go \[\sin x(1-2\cos x)=0\]

OpenStudy (xapproachesinfinity):

set EACH ONE EQUAL TO ZERO

OpenStudy (xapproachesinfinity):

and solve

OpenStudy (xapproachesinfinity):

let me see your work!

OpenStudy (anonymous):

how did you get sin x (1-2cos x)=0?

OpenStudy (xapproachesinfinity):

i factored sinx my friend

OpenStudy (anonymous):

wait did you subtract 2sin x from sin x?

OpenStudy (xapproachesinfinity):

no do you know how to factor?

OpenStudy (anonymous):

Yeah I can factor, but I get confused with trigonometric functions

OpenStudy (xapproachesinfinity):

how do you factor \[x-2xy\]

OpenStudy (xapproachesinfinity):

say how would you factor?

OpenStudy (anonymous):

I have no idea

OpenStudy (xapproachesinfinity):

you just said you know how to do it, but you only have issue with trig

OpenStudy (xapproachesinfinity):

that's not trig just a linear expression

OpenStudy (anonymous):

I can factor equations like x^2+3x+3

OpenStudy (xapproachesinfinity):

well if you can factor that why can't you factor x-2xy? what is common btw x and -2xy?

OpenStudy (anonymous):

the x's

OpenStudy (xapproachesinfinity):

yes so can we then factor the x's and get x(1-2y)

OpenStudy (xapproachesinfinity):

if i factored x in the first i'm left with 1 if i factored x in the second term i'm left with -2y

OpenStudy (anonymous):

sorry my net cut out

OpenStudy (anonymous):

sin x=0 cosx= 1/2

OpenStudy (anonymous):

0, pi, pi/6, 5pi/6 is what I got

OpenStudy (xapproachesinfinity):

show your work?

OpenStudy (anonymous):

I just used the unit circle?

OpenStudy (xapproachesinfinity):

show the whole work my friend, you skipped many steps

OpenStudy (anonymous):

sin x=0 1-2cosx=0 -2cosx=-1 cox x=1/2 ?

OpenStudy (xapproachesinfinity):

yes and then?

OpenStudy (anonymous):

I just used the unit circle, sinx=0 is at pi and 0 and cosx=1/2 is at pi/3 and 5pi/3

OpenStudy (anonymous):

damn I messed it up lol

OpenStudy (xapproachesinfinity):

what about 2pi?

OpenStudy (xapproachesinfinity):

for sin

OpenStudy (xapproachesinfinity):

isn't zero as well

OpenStudy (xapproachesinfinity):

and it is pi/6 not pi/3 the first answer you provided was correct except that you missed 2pi

OpenStudy (anonymous):

Yeah I got 0, 2pi I'm assuming because it's [0,2pi)?

OpenStudy (xapproachesinfinity):

oh it is open i didn't see that then you are good

OpenStudy (anonymous):

alright thanks, can you help me with one more?

OpenStudy (anonymous):

sin(2x+pi/6)=1/2 no idea how to even begin

OpenStudy (xapproachesinfinity):

when does an angles give 1/2 when you sin it? and is there and interval given (because you have to make sure the values are in that interval)

OpenStudy (anonymous):

interval is [0,2pi) and pi/6 and 5pi/6

OpenStudy (xapproachesinfinity):

then 2x+pi/6=pi/6 or 2x+pi/6=5pi/6

OpenStudy (xapproachesinfinity):

solve for x

OpenStudy (anonymous):

x=2 or x=4pi/3 ?

OpenStudy (xapproachesinfinity):

how come x=2? 2x=0 how did you get x=2?

OpenStudy (anonymous):

I read it wrong, oops

OpenStudy (anonymous):

so sin of 0 and sin of 4pi/3?

OpenStudy (xapproachesinfinity):

hmm no x=0 and x=4pi/6=2pi/3

OpenStudy (anonymous):

I thought 4pi was its own entity? I'm doing online math so I'm essentially teaching everything to myself so sorry for the lack of math experience

OpenStudy (xapproachesinfinity):

oh no problem you did good you just didn't simplify nicely 5pi/6-pi/6 =4pi/6 not 4pi/3

OpenStudy (anonymous):

so the next step would be to get the sin of both angles?

OpenStudy (anonymous):

or the sin of 1/2?

OpenStudy (xapproachesinfinity):

you are done what else you want? they asked you to find x and you found it??

OpenStudy (anonymous):

why is the answer 0,pi/3,pi,4pi/3? I thought it was 2pi/3?

OpenStudy (anonymous):

where did the pi/3 and 4pi/3 come from?